OpenStudy (anonymous):

are the two lines parallel x+3y=-4,3x+9y=8

OpenStudy (anonymous):

parallel lines must have the same slope

OpenStudy (anonymous):

you can divide the second equation by three to compare them

OpenStudy (anonymous):

So they are not parrallel

OpenStudy (anonymous):

I am solving it now and see

OpenStudy (anonymous):

one of them has 3y and the other has 9y yes

OpenStudy (anonymous):

you can't compare them directly

OpenStudy (anonymous):

Oh I solved it they are not parallel

OpenStudy (anonymous):

Divide the second equation by 3

OpenStudy (anonymous):

are you sure ?

OpenStudy (anonymous):

There are Parallel

OpenStudy (anonymous):

Yes I am sure accordint to my working ,You too can do the work and see the final answer ,they are not parallel

OpenStudy (anonymous):

@ Yahoo How did you think they are parallel because I go against it do let me know

OpenStudy (anonymous):

x+3y=-4, 3x+9y=8 first one: x+3y=-4 3y = -4 -x y = -4/3 -1/3x second one: 3x+9y=8 9y = 8-3x y = 8/9 - 3/9x y=8/9 -1/3 x

OpenStudy (08surya):

yes they are parallel becoz both have slope=-1/3

OpenStudy (anonymous):

Well parallel lines have the same gradient(slope) get those in the form y=mx+c where m is the gradient and check it wether both the gradients are equal if they are they are parallel

OpenStudy (anonymous):

I am checking it on now in order to agree with you guys that they are parallel

OpenStudy (texaschic101):

they are parallel x + 3y = -4 3y = -x - 4 y = -1/3x -4/3 <-- slope of -1/3 3x + 9y = 8 9y = -3x + 8 y = -3/9x + 8 y = -1/3x + 8 <-- slope -1/3

OpenStudy (anonymous):

I agree with you guys